Write all the factors of the following numbers: 
18 ????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

18 = 1 × 18 18 = 2 × 9 18 =3 ×6 18 = 6 × 3
Stop here since 3 and 6 have occurred earlier
Hence, the factors of 18 are 1, 2, 3, 6, 9 and 18

Thanks